OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (2,-3) and is perpendicular to x-2y=5

OpenStudy (anonymous):

Now what do you know about perpendicular lines? esp. their slopes?

OpenStudy (anonymous):

They have 90 degree angles

OpenStudy (anonymous):

and are negative reciprocals of each other

OpenStudy (anonymous):

right! well, the slope of x-2y=5 is -a/b, which is -(1/-2)=1/2 so we would take the negative reciprocal of 1/2 which is -2. -2(2)+b=-3 -4+b=-3 b=3/4 which is our y-int. so our equation is y= -2x+3/4

OpenStudy (anonymous):

Thanks, my answer key states the right answer is y= -2x+1?

OpenStudy (anonymous):

oh whoops. yes. from -4+b=-3 I accidently divided. it is -3+4=1 so indeed it is -2x+1. sorry about that

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

Where did you get the positive 2 that you multiplied from the -2?

OpenStudy (anonymous):

you take the given coordinate (2,-3) and substitute into the x and of: -2x+b=y then you solve for b. (2,-3) is a part of the line, and we know it should have a slope of -2. so now what we had left to do was solve for b, or the y-intercept. Then once we have b and the slope, it can become a function, ready to accept any given x-value to put out a yvalue. these values are on the line

OpenStudy (anonymous):

Thank you! I've got it now. Can't thank you enough!

OpenStudy (anonymous):

no problem! I really recommend Khan academy's videos, or buying some Art Of Problem Solving workbooks. they're both very helpful.